Given that M =
2 −1
−3 4
and that M2 − 6M + kI = 0, find k
(Under matrix )
1
Expert's answer
2020-05-28T19:40:14-0400
"M=\begin{pmatrix}
 2 & -1 \\
 -3 & 4
\end{pmatrix}"

"M^2=\begin{pmatrix}
 2 & -1 \\
 -3 & 4
\end{pmatrix}\begin{pmatrix}
 2 & -1 \\
 -3 & 4
\end{pmatrix}="

"=\begin{pmatrix}
 2(2)-1(-3) & 2(-1)-1(4) \\
 -3(2)+4(-3) & -3(-1)+4(4)
\end{pmatrix}="

"=\begin{pmatrix}
 7 & -6 \\
 -18 & 19
\end{pmatrix}"

"-6M=-6\begin{pmatrix}
 2 & -1 \\
 -3 & 4
\end{pmatrix}="

"=\begin{pmatrix}
 -6(2) & -6(-1) \\
 -6(-3) & -6(4)
\end{pmatrix}=\begin{pmatrix}
 -12 & 6 \\
 18 & -24
\end{pmatrix}"

"M^2-6M=\begin{pmatrix}
 7 & -6 \\
 -18 & 19
\end{pmatrix}+\begin{pmatrix}
 -12 & 6 \\
 18 & -24
\end{pmatrix}="

"=\begin{pmatrix}
 -5 & 0 \\
 0 & -5
\end{pmatrix}=-5\begin{pmatrix}
 1 & 0 \\
 0 & 1
\end{pmatrix}=-5I"

"M^2-6M+kI=-5I+kI=0"

"k=5"